Liquid octane ( C 8 H 18 C8H18 ) has a density of 0.7025 0.7025 g/mL at 20 °C. Find the true mass ( ? true mtrue ) of octane when the mass weighed in air is 18.800 g. Assume the air density is 0.0012 g/mL and the balance weight density is 7.7 g/mL.
Answer #1

Solution:

The true mass (Wt) is calculated as,

Wt = Wa x [ 1 + (1/Do - 1/Dw) x Da]

Where, Wa = Weight in air = 18.800 g

Do = Density of object = 0.7025 g /mL

Dw = Density of balance weight = 7.7 g /mL

Da.= Density of air = 0.0012 g /mL

On putting the above values, we get

Wt = 18.800 [1 + (1/0.7025 - 1/ 7.7) x 0.0012]

Wt = 18.800 [1 + (1.4235 - 0.1298) x 0.0012 ]

Wt = 18.800 [ 1 + (1.2937 )x 0.0012]

Wt = 18.800 + 18.800 x 0.00155

Wt = 18.800 + 0.0292 = 18.8292 g